OpenStudy (anonymous):

intergrate cos^2 (3x) dx

OpenStudy (anonymous):

rewrite it as \[\frac{1}{2}\cos(6x)+1\]

OpenStudy (anonymous):

Thanks, i appreciate it.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wow i made a typo sorry

OpenStudy (anonymous):

forgot parentheses, it should be \[\frac{1}{2}\left(\cos(6x)+1\right)\]

OpenStudy (anonymous):

using the identity \[\cos^2(x)=\frac{1}{2}\left(\cos(2x)+1\right)\] only with \(3x\) instead of \(x\)

OpenStudy (anonymous):

you do sine the same way

OpenStudy (anonymous):

don't worry i understood what you meant

OpenStudy (anonymous):

k good